Ten kilograms of water enters the heater at 16 0 C and leaves at 123 0 F. Determine

the heat added in KJ.


1
Expert's answer
2021-10-13T15:43:38-0400

"Q=cm(t_2-t_1)=\\=4200\cdot 10\cdot(50.6-16.0)=\\=1453~\text{kJ}."